OpenStudy (cookiemonster18):

I'll fan and medal :) Simplify the radical expression. √363-3√27 (Show all steps)

OpenStudy (andrewyates):

\[\sqrt{363} = \sqrt{121*3} = 11\sqrt{3}\] and \[3\sqrt{27} = 3\sqrt{9*3} = 9\sqrt{3}\] Puting this together: \[\sqrt{363} - 3\sqrt{27} = 11\sqrt{3} - 9\sqrt{3} = 2\sqrt{3}\]

OpenStudy (anonymous):

^ What he said.